# Emotional Strain and Frustration Drivers in LLM-Human Interactions in Software Engineering

### v1

@@ IV-A Respondents Demographics
The responses in our survey showed a diverse sample of software engineers in terms of ages (See Figure 3), and roles (See Table I). The organizations where across 7 countries distributed over 3 continents and are within the domains: Airline and Aircrafts, Automotive, Game design, Infotainment, eLearning, Cybersecurity, Telecommunication, Trade, and Software Engineering research and education.
Figure 3: The distribution of the ages of the participants.
TABLE I: Participants’ areas and roles.
@@
Regarding the nature of their familiarity with LLMs, the vast majority of our participants (58 out of 62) describe themselves as “familiar” or “very familiar” with LLMs. Our participants also describe the wide range of LLMs that they use at work in Figure 4 with ChatGPT being the most popular one.
IV-B Usage of LLMs in Software Engineering industry and academia
The results show that 56 participants (94.9%) use LLMs occasionally, out of which 38 participants (66%) use them on a weekly or daily basis. Only 3 (5.1%) participants indicated that they rarely use LLMs at work. Regarding the LLMs that the participants use at work, the results (depicted in Figure 4) show that ChatGPT was by far the most used one. Table II shows the tasks for which academia and industry respondents apply them. We considered more than once answer per participant.
Figure 4: LLMs used by our participants.
In industry, respondents use LLMs for programming tasks like code generation, debugging, and optimisation. They also employ LLMs for written communication tasks, such as drafting emails and brainstorming ideas, and for generating and improving text. Additionally, LLMs help users onlearning new technologies and research by providing starting points, best practices, and summaries of lengthy information. Lastly, respondents view LLMs as digital assistants for task management and problem-solving, streamlining workflows and enhancing productivity.
TABLE II: LLM Usage Across Industrial and Academic Tasks
@@ IV-B Usage of LLMs in Software Engineering industry and academia
IV-C Emotions during LLM interaction
We analyse the emotional responses described by our participants when interacting with LLMs during work, more specifically when the interaction takes a different path than expected and provide an incorrect response. Since emotional responses are complex and layered rather than straightforward, there was a one-to-many relationship between participants and emotions, where participants described a combination of multiple emotions when describing their experiences.
We used the emotions wheel by Wilcox (See Figure 1) to map participants’ answers to various emotional states, allowing us to categorise and analyse the range of emotions experienced when using LLMs. According to Willcox’s emotions wheel, frustration comes initially from anger, which can develop into annoyance, then eventually frustration. These three emotions make up more than half of the emotions observed among our participants in Figure 5. More specifically, the most common feelings, that is, frustration (30.4%) and annoyance (17.4%), can challenge software engineers’ well-being, and a continuous exposure to these emotions can also create a barrier to smooth workflows and a cumulative risk of emotional strain over time.
Other participants reported feeling indifferent (13%), while some showed disappointment (10.1%) which are related to how the engineers’ expectations are met by the LLM.
@@ IV-C Emotions during LLM interaction
Participant 29: “Knowing how LLM[s] work, I typically have lower expectations. So I [don’t] feel as frustrated or disappointed, particularly if I know that the task I asked is not trivial. I just try again in a different way.”
Such expectations come from building knowledge about the LLM being used as well as understanding its capabilities and limitations based on previous interactions.
Less frequent reactions included positive emotions like calmness (5.8%), playfulness (2.9%) and curiosity (1.4%), as well as negative emotions such as anger (2.9%), confusion (2.9%), stress (1.4%), and sadness (1.4%). Some participants also reported feeling thoughtful (1.4%) or even guilty (1.4%) showing the varied and sometimes unexpected emotional responses that emerged. These emotions suggest that interacting with LLMs is not just a functional exchange but an exploratory experience for some software engineers. Curiosity, for example, has attached an investigative mindset often aligned with a trial-and-error approach. Playfulness, meanwhile, shows a willingness to engage with the LLM on a more relaxed, open-ended basis. Guilt, though rare, introduces a new angle, suggesting that some engineers may feel a sense of responsibility or even self-blame if the interaction does not go as expected.
Figure 5: Emotional responses when receiving an incorrect answer. The colors map to Willcox’s emotions wheel in Figure 1.
IV-D Expectations when interacting with LLMs
As shown in Table III, participants’ expectations when using LLMs go beyond mere functionality and extend into deeper aspects of quality, reliability, and versatility. For quality and performance, software engineers expect LLMs to consistently deliver correct and unhallucinated information without error or delay, as inconsistency can erode trust. The participants also indicated that LLMs must be user-friendly, that is, ensure a smooth and easy interaction with the user while ensuring data protection, reflecting participants’ concerns over privacy in sensitive tasks. In terms of understanding, engineers value LLMs that grasp context and intent, highlighting the need for deeper comprehension of nuanced queries. Some added that they expect the LLM to understand the context without the need for an excessive description of the context in the initial prompt, but also ask clarification questions if needed to better understand the intent and the context.
@@ IV-D Expectations when interacting with LLMs
Participant 38: “I expect LLMs to recognize and alert me if they are lacking certain information from my prompt”
When it comes to answer structure, most of the participants who use LLMs for research or formatting text emphasised the importance of clear, organised, and grammatically correct responses. In addition, engineers preferred concise or elaborative answers depending on their needs and the task they want support with. For code-related tasks, engineers preferred that the response starts with the snippet of code that is needed (e.g., to repair a program) followed by an explanation that the engineer can optionally go through. While elaborative answers where perceived beneficial for engineers who used LLMs to learn new concepts or explain certain artifacts (e.g., error message). Information integrity was also expected from LLMs. More specifically, the participants mentioned how the LLM should be able to provide the sources of information that were used to generate the output, as well as the confidence estimation of the accuracy of the output. The participants motivated these two transparency criteria as a means to enable the verification of the information and support decision making (e.g., deciding how to utilize the output). Finally, in terms of versatility, our findings showed that integration with other tools and adaptability to different workflows are increasingly expected by engineers in industry as the adoption of LLMs is increasing in software organisations. This adaptability allows the LLMs to fit into varied contexts and communication preferences, signalling a strong need for flexibility.
TABLE III: Users’ Expectations when Using LLMs
@@
Balance between conciseness and thoroughness (14)
@@
Domain and Contextual Understanding (2)
@@
IV-E Frustration triggers in software engineering
We ask our participants to describe one or more situations of where they felt frustrated when interacting with an LLM. Regardless of the area of Software Engineering (industry or academia), we observed patterns of triggers that cause the LLM to fail to meet the software engineer’s expectations and hence cause frustration. In the following, we describe the triggers of a frustrating interaction between software engineers and LLMs.
Incorrect answers and hallucinations: One of the most common causes of frustration was continuously receiving incorrect or hallucinated responses from the LLM. While receiving such responses was referred to as “Annoying” and “Disappointing”, the frustration is caused when these responses are received continuously even when the engineers change their prompt or try to correct the LLM. Our participants described such situations as LLMs being stubborn and insisting on the an incorrect or hallucinated answer.
@@ IV-E Frustration triggers in software engineering
The definition of incorrectness varied among participants. Some examples of incorrect answers are uncompilable or buggy code, incorrect explanation of error message or incorrect factual information that were verified using other sources (e.g., an expert or a search engine). Hallucination was described as nonsense explanations, refernces that do not exist, functions and syntax in a code that do not exist.
@@
Intent not understood: Frustrations (or emotions that can evolve to frustration) are also caused when the software engineers get the impression that the LLM did not understand their prompt. Not understanding can be reflected on the response with an irrelevant response to the initial question.
@@
Participant 8: “I asked CodeLlama to explain the code of a function in a library with many lines, and it proceeded to explain a different function from a different library.”
However, a few of the participants pointed out that such misunderstandings can be caused by the poor phrasing or the lack of context in the prompt.
@@
Personal preferences unmet: Many of our participants pointed out that they get frustrated with reasons that are related to their personal preferences. Some aspects of certain LLMs (for example, the answer structure) can be annoying to some engineers, and when the frequency of interactions with the LLM increases, the annoyance turns into frustration. For example, two of participants pointed out that the way that ChatGPT apologized every time they tried to correct the LLM or send a follow-up prompt was a source of frustration. Others were frustrated with how the LLM they use only provides answers that are long, or only in a bullet points. Forcing the LLM to structure an answer that aligns with their preferences required specifying many requirements and constraints in the prompt.
@@
Participant 35: “[I get frustrated] when ChatGPT gives to long answers. I quite often ask things that can be answered with a short sentence, but still I get half a page of ramblings back.”
Limitations of the LLM: Other frustration triggers were due to limitations of the LLM. Such limitations can lead the LLM to be unable to perform a specific task if it is not trained on to do so, or could be a limitation with the configurations of the LLM e.g., the size of the context window. Since software engineers cannot do anything to overcome these limitations, they get frustrated when they attempt to force the LLM to overcome limitations by prompting. For instance, a software developer can become frustrated when attempting to create a complex and large application using one prompt, but the LLM stops generating the response when the maximum output token is reached [^1] https://huggingface.co/docs/transformers/en/llm_tutorial resulting in an incomplete application code with missing lines of code with logic and syntax errors.
In these cases, the software engineer can use other LLMs that overcome these limitations rather than trying to prompt the same LLM differently.
@@
After describing their experiences, the participants indicated their perceived importance level of five aspects of an LLM, that is, Correctness, Hallucination, Understanding, Performance, and Ability to Answer, which can impact their overall experience. We report the importance levels that followed a 5-level Likert Scale in Figure 6. We found that the importance levels align with the frustration triggers we identified. Correctness, hallucination and understanding were assigned higher importance levels, indicating their high impact on the engineer’s experience when interacting with LLMs. Whereas the performance of the LLM (e.g., response time), and the ability to answer a question where deemed as less important.
Figure 6: Likert-scale results of the importance level of different aspects of the LLMs that can impact the user experience.
IV-F Unmet expectations’ impact on motivation
When the LLM fails to assist our participants, the participants’ motivation to complete their task was influenced mainly in three ways.
21.3% of our respondents reported that their motivation decreased when LLMs did not give the correct answer. Responses expressed frustration, stress, or disappointment, impacting participants’ willingness to continue. “My motivation decreases. I try to give it more context or constraints to perform the task”.
Some other participants commented that they eventually gave up. These respondents mentioned that after some effort, they decided to abandon the LLM and move on to other methods or stop altogether.
Another group was formed of 77% of the participants who were unfazed by the LLM’s failure, treating it as a non-critical tool and continuing with the task with their motivation not being affected.
@@ IV-F Unmet expectations’ impact on motivation
In an interesting case, one participant mentioned that their motivation increased when the LLM did not assist them as intended, this was due to a desire to solve the problem manually or a learning experience.
@@
IV-G Suggestions and ways to reduce frustration
When asked what actions participants typically take after receiving an incorrect answer from an LLM, the majority (41 out of 62 participants) said they changed the prompt to try again. A smaller group (12 participants) reported providing feedback to the LLM, while a few (2 participants) said they did nothing. See Figure 7.
@@ IV-G Suggestions and ways to reduce frustration
In the “Other” category, participants described various strategies. Some combine multiple actions, such as changing the prompt, switching to another LLM, or even reverting to traditional search methods like Google or Wikipedia. A few participants mentioned disengaging from the LLM entirely or writing the response themselves.
@@
Figure 7: Actions done after receiving a wrong answer from LLMs
@@
We asked our participant about the improvements they would like to see that would enhance their experience when using LLMs. They identified several areas; we grouped them and show the results in Table IV. The suggestions mainly align with the engineers’ expectation (See Table III). Regarding the quality of the outcome, the participants emphasized the importance of reducing hallucinations and improving response accuracy. They also commented on the need for enhanced analytical capacity to tackle more complex problems.
Regarding usability, participants expressed a preference for concise and short answers (unless they are prompted otherwise) and user-friendly interaction. They also stressed the importance of memory and learning capabilities, allowing LLMs to retain relevant information and build on past interactions. For understanding, prompt comprehension was considered essential, with users expecting LLMs to ask clarifying questions when necessary. Additionally, participants wanted transparent communication about the system’s limitations.
When it comes to the answer structure, users called for consistency in responses and better organisation of information. They also noted the need for LLMs to handle simple formatting tasks more efficiently. As for information quality, participants focused on the importance of correct references and proper citation of sources. They also mentioned concerns about confidentiality, urging transparency in how sources are used and shared. Finally, regarding versatility, users indicated the importance of handling more complex tasks and adapting to different user needs. Increased functionality to support a broader range of tasks was also crucial for improving overall user satisfaction.
TABLE IV: Improvements for Better User Experience
@@
Performance Higher memory utilization (4)
@@
V Discussion
V-A Frustrations in the context of LLM interaction
We identified four main triggers that cause frustrations or feelings that can develop into frustration over the time. These causes are related to correctness and hallucinations, misunderstandings, personal preferences, and limitations of the used LLM. In code-related tasks, correctness and meeting personal preferences were important for the software engineers, since the absence of either can make the response from the LLM unusable. For instance, if the code is buggy or does not follow a standard practice within the software organization, it creates additional work. When the outcome is a piece of code and it is buggy, it requires the developer to perform debugging rather than implementation, which can be expensive in terms of time when the developer is working on code not written by them [fitzgerald2008debugging], especially, if it is written by an LLM which can produce less maintainable code [liu2024refining] and can be more frustrating than writing the code with no LLM assistance [weisz2022better]. In other cases, when the LLM is unable to provide a response due to a limitation of the model or a misunderstanding of the engineer’s intent (especially when the prompts is specific software-related research area), it can cause disruption to the workflow of the engineer when they attempt to make the LLM understand and provide the expected outcome, as we see in Figure 7, where most of the participants provide feedback, change the prompt, or change the LLM when they encounter an incorrect response. As a consequence, the progress of the software engineer can be delayed, making them feel as though they have spent longer than planned to reach their goal (e.g., completing a task), which enforces frustration and emotional strain [ceaparu2004determining, lazar2006workplace].
Furthermore, frustrating experiences among our participants resulted from several emotions other than frustration directly, such as anger and annoyance (See Figure 5). Such emotions discussed in studies that focused specifically on a GitHub Copilot [^2] https://github.com/features/copilot, an LLM-powered code assistant. Eshraghian at al. [eshraghian2024ai] shows that emotions such as frustration and anger can come from the feeling of a threat without being able to control it. For instance, practitioners where annoyed from the data privacy policy of GitHub Copilot, which has a threat of leaking confidential data (e.g., production code), with very little control over it (i.e., to use the LLM, they need to accept the policy). Frustrations caused by GitHub Copilot also centred around usability [usability2023frustration], where providing code suggestions, especially large ones, when they are not needed made developers frustrated. These Copilot-specific frustration causes were not observed specifically among our participants since very few of them used GitHub Copilot (See Figure 4). However, providing too large code suggestions and the need for the LLM to use the code data fit under our defined frustration triggers “failing to meet personal preferences” and “limitations of the LLM” respectively.
We distinguish the frustration triggers that we identified from frustrations that can be caused by technologies in other domains. In the general domain, frustrations caused by tools and technologies are due to performance issues (e.g., system crash) [lazar2006workplace], which we did not observe among our participants. While our participants mentioned that they expect the LLMs to be stable and efficient in terms of speed of response (See Table III), they were not frustrated because of any performance issues. We argue that this is due to recent and popular LLMs being more stable and efficient. Furthermore, technologies adopted in the medical domain are usually heavily relied on for clerical tasks (e.g., documentation), and frustrations can occur when the user interface is complex [opoku2015user], but are also associated with the nature of emotionally exhausting the work environment, which can enable frustration when working with technology at work [tawfik2021frustration]. We observed among our participants that LLMs are not essential for them to complete their tasks as they can rely on their expertise or use other alternatives (See Section IV-F). While in other domains, technologies can be more crucial and relied on e.g., using a technology to document a patient’s state.
Takeaway: The four main frustration triggers come from potential things that can require extra processing time (e.g., modifying an output) and can disrupt the workflow of the engineer.
V-B Impact of frustrating experiences on motivation
While most of our participants expressed frustration (or similar emotions), their motivation to complete the task was not necessarily affected. Our results showed that participants often felt demotivated when an LLM failed to meet their expectations, such as when it did not assist them as intended, frequently leading to frustration. However, most of those who felt frustrated reported that their motivation remained intact, likely due to perceiving the LLM as merely one tool among others in achieving their goal. When an LLM could not provide the necessary assistance, participants commonly resorted to alternative solutions, such as using search engines like Google or relying on their expertise to complete the task independently. These observations align with findings by Franca et al. [francca2014motivated], who explored the connection between motivation and the satisfaction and happiness of software engineers, revealing that happiness slightly overlaps with but does not correlate with motivation.
These findings suggest that, despite facing emotional challenges, software engineers maintain their motivation to persist with demanding tasks.

### v2

@@ 4.1. Respondents Demographics
Our survey sample included software engineers in diverse roles (see Table 1) with a median age of 32. Participants represented organisations from seven countries across three continents, spanning aviation, automotive, game design, infotainment, eLearning, cybersecurity, telecommunications, trade, and SE research and education domains.
Table 1. Participants’ areas and roles.
@@
Most participants (58 of 62) described themselves as ‘familiar” or “very familiar” with LLMs. Figure 3 shows the range of LLMs they use at work, with ChatGPT being the most popular.
4.2. Usage of LLMs in Software Engineering Industry and Academia
The results show that 56 participants (94.9%) use LLMs occasionally, out of which 38 participants (66%) use them on a weekly or daily basis. Only 3 (5.1%) participants indicated that they rarely use LLMs at work. Table 2 shows the tasks for which academia and industry respondents apply them. We considered more than one answer per participant.
Figure 3. LLMs used by our participants.
[…]
[LLMs]LLMs used by our participants.
In industry, respondents use LLMs for programming tasks like code generation, debugging, and optimisation. They also employ LLMs for creative and communication tasks, such as drafting emails and brainstorming ideas, and for generating and improving text. Additionally, LLMs help users on learning new technologies and research by providing starting points, best practices, and summaries of lengthy information. Lastly, respondents view LLMs as digital assistants for task management and problem-solving, streamlining workflows and enhancing productivity.
Table 2. LLM Usage Across Industrial and Academic Tasks
@@ 4.2. Usage of LLMs in Software Engineering Industry and Academia
4.3. Emotions During LLM Interaction
Due to the complexity of the emotional responses to unexpected LLM interactions, participants often described multiple, layered feelings in their experiences. We mapped these feelings described to Willcox’s emotions wheel (see Figure 1) to assess on a more fine-grained level how evolved the emotions were. Following this framework, we could trace how initial feelings, such as anger (first-level emotion), might progress into more specific emotions, like annoyance (second level) and then frustration (third level), and quantify how often these emotions occurred at each stage.
In Figure 4, we show the different emotions that were expressed by our participants for each category that map to Willcox’s emotions wheel. For instance, we found that the most common emotions (54.6%) are frustration or emotions that can develop into frustration, such as annoyance and anger. This poses potential challenges to well-being and smooth workflows, with a risk of cumulative emotional strain over time.
Many respondents (27.8%) have also reported sadness-related emotions such as disappointment, indifference, or even guilt. In contrast to anger-related emotions, where respondents primarily blamed the LLM for its limitations, those who felt disappointed, sad, or guilty often turned the blame inward and criticised themselves for not being able to write the right prompt or meet their own expectations. When expectations were lower, the disappointment turned into indifference.
@@ 4.3. Emotions During LLM Interaction
Participant 29: “Knowing how LLM[s] work, I typically have lower expectations. So I [don’t] feel as frustrated or disappointed, particularly if I know that the task I asked is not trivial.”
Such expectations come from building knowledge about the LLM and understanding its capabilities and limitations based on previous interactions.
Less frequent reactions included positive emotions like calmness, thoughtfulness, playfulness, and curiosity, as well as negative emotions such as confusion and fear. This shows the varied and sometimes unexpected emotional responses that emerged. These emotions suggest that interacting with LLMs is not just a functional exchange but an exploratory experience for some software engineers. Curiosity, for example, has attached an investigative mindset often aligned with a trial-and-error approach. Playfulness, meanwhile, shows a willingness to engage with the LLM on a more open-ended basis. Fear introduces a new angle, suggesting that some engineers may feel a sense of responsibility if the interaction does not go as expected.
Figure 4. Emotional responses when receiving an incorrect answer. The colors map to Willcox’s emotions wheel in Figure 1.
[…]
4.4. Expectations When Interacting with LLMs
As shown in Table 3, participants’ expectations for LLMs extend beyond functionality to quality, usability, and versatility which are important factors in developing effective and trustworthy products in software development and design. Regarding quality and performance, engineers expect LLMs to consistently deliver correct and unhallucinated information without error or delay, as inconsistency can erode trust in the development workflow.
@@ 4.4. Expectations When Interacting with LLMs
Participant 19: “I rely on the LLM to provide accurate, relevant information that I can trust for both coding tasks and daily life management. It’s like having an expert who gets it right the first time!”
In terms of understanding, engineers value that LLMs understand context and intent, preferring models that ask clarification questions when necessary. They expect the LLM to understand the context without the need for a detailed context description in the initial prompt. This expectation is important since software research and development happen in dynamic and complex environments that require a lot of context such as best practices, policies, and relevant software artifacts. With the lack of consideration for such contexts, the outcome can become unusable and hard to integrate into the solution.
[…]
Participant 19: “I need an LLM that can seamlessly adapt to different contexts. Whether it’s helping me with technical jargon, understanding project management lingo, […], the LLM should be versatile enough to handle it all.”
Furthermore, since software engineering is a broad field with researchers from various sub-domains and specialties, researchers need to organize their texts and use language that fits the targeted audience. Therefore, software researchers both in academic and industrial organizations emphasized the importance of clarity and organisation of LLM-generated text, and that tailoring the answer to the structure that the task needs are crucial. For code-related tasks, engineers prefer responses starting with code snippets followed by explanations. When using LLMs to learn new concepts or explain artefacts, participants preferred elaborative answers.
Additionally, they expect LLMs to be transparent by providing the source of the information and confidence estimation of the output accuracy to ensure information integrity and support decision-making throughout the development lifecycle. This should also come with a need to protect the shared information in the chat.
Finally, in terms of versatility, engineers in industry increasingly expect LLMs to integrate with other tools and adapt to diverse workflows, reflecting the growing need for flexibility in software engineering environments. This also requires a usable LLM, ideally with a user-friendly interface that enables intuitive interactions and a seamless integration of the LLM in the software-related tasks.
Table 3. Users’ Expectations when Using LLMs
@@
Conciseness vs. thoroughness (14)
@@
Domain/Contextual Understanding (2)
@@
4.5. Frustration Triggers in Software Engineering
After exploring emotions in general related to receiving unexpected answers from LLMs, we focus on frustration-related emotions. We asked participants to describe specific situations where they felt frustrated when interacting with LLMs. From this, we identified patterns of triggers that cause LLMs to fail to meet engineers’ expectations, leading to strains such as frustration. We outline the frustration triggers below.
Repeated inaccuracies and hallucinations: One of the most common causes of frustration was receiving repeated incorrect or hallucinated responses from the LLM.
The definition of incorrectness varied among participants. Some examples of incorrect answers were uncompilable or buggy code, incorrect explanations of error messages, or incorrect factual information that was verified using other sources (e.g., an expert or a search engine). Hallucination was described as nonsense explanations, references that do not exist, made-up packages, and invalid syntax in a programming language.
Participants explicitly described that frustration arose when these issues persisted despite attempts to rephrase prompts or correct the LLM. For instance, participants referred to such situations as ”annoying” or ”disappointing” initially. However, they noted that repeated failures led to frustration, describing the LLM as “stubborn” and “insisting on an incorrect or hallucinated answer”.
@@ 4.5. Frustration Triggers in Software Engineering
This pattern of repeated failures can disrupt workflows in software engineering, where development cycles are often fast-paced and agile which requires more reliable and stable tools. For example, when the LLM provides a code that imports hallucinated libraries renders the code unusable which leads to wasted time fixing, debugging, or rewriting the entire implementation.
@@
Intent not understood: Frustrations (or related emotions) are also caused when the software engineers feel that the LLM did not understand their prompt. Not understanding can be reflected in the response that is irrelevant to the initial question. Intent understanding was a common frustration trigger among our participants from industry and academia since in practice, engineers’ queries are often highly technical and domain-specific and deal with complex software artefacts. Similarly, researchers and academics deal with novel techniques and niche problems that may cause the LLM to misunderstand the intent. Note that such misunderstandings are more common in general-purpose LLMs than in fine-tuned and specialized LLMs.
@@
Participant 29: “I was trying to ask [LLM] to fill one specific cell in a notebook based on the others but it kept returning the same generic code for two cells instead of one. I had to talk to [LLM] like a child and say don’t do that and do this, and only this.”
On another note, some participants pointed out their perceived usefulness of prompt programming and carefully constructing a prompt that would minimise such misunderstandings that can be caused by poor phrasing or the lack of context in the prompt.
@@
While prompt programming has shown a high potential in enhancing the LLM outcome, it remains unclear whether it is effective in software engineering-related tasks.
Personal preferences unmet: Many of our participants pointed out that they get frustrated with reasons related to their personal preferences. Some aspects of certain LLMs (e.g. answer structure) can be annoying to some engineers, and when the frequency of interactions with the LLM increases, the annoyance turns into frustration. For example, two participants pointed out that an LLM apologising every time they tried to correct the LLM was a source of frustration since it can disrupt the workflow, especially during a refactoring or debugging process with the LLM, which results in a long conversation with many follow-up prompts. Others were frustrated with how the LLM they use only provides answers that are long, or only in bullet points. Forcing the LLM to structure an answer that aligns with their preferences required specifying many requirements and constraints in the prompt.
@@
Participant 35: “[I get frustrated] when [LLM] gives too long answers. I quite often ask things that can be answered with a short sentence, but still I get half a page of ramblings back.”
Such preferences depend on the task that the software engineer is solving. For example, important details when debugging might be hidden in long responses, while overly concise bullet points might omit crucial information needed to understand a system’s architecture.
Limitations of the LLM: LLM limitations (e.g., inability to perform specific tasks) or configuration constraints (e.g., context window size) are frustration triggers for software engineers. When an engineer attempts to force the LLM to overcome these limitations by prompting, it often leads to more frustration. For example, trying to generate a large application code in one prompt can result in missing lines and errors when the LLM hits its maximum token limit (HuggingFace, 2023). The participants highlighted that there are some of the many tasks in software engineering research and practice that LLMs are just “not good at”.
@@
Finally, as a verification question, we asked the participants to rate aspects of LLMs (correctness, lack of hallucinations, understanding, performance, and ability to answer) on a 5-point Likert scale based on how important they are in an LLM in order to ensure a better user experience. The results in Figure 5 resonate with the previous results where correctness, lack of hallucination, and correctness being the most important. In comparison, performance (e.g., response time) and ability to answer were seen as less critical.
Figure 5. Likert-scale results of the importance level of different aspects of the LLMs that can impact the user experience. The scale ranges from Not Important (left - red) to Extremely Important (right - green).
[…]
4.6. Unmet Expectations’ Impact on Motivation
When the LLM failed to assist our participants, the participants’ motivation to complete their task was influenced mainly in three ways.
[…]
Participant 29: “Eventually I give up [on the task], or report negative experimental results.”
Another group was formed of 77% (48 out of 62) of the participants who were unfazed by the LLM’s failure, treating it as a non-critical tool and continuing with the task with their motivation not being affected.
@@ 4.6. Unmet Expectations’ Impact on Motivation
In an interesting case, one participant mentioned that their motivation increased which was due perceiving the interaction as a learning opportunity.
@@
4.7. User Actions for Improving LLM Interactions
@@
When asked what actions participants typically take after receiving an unexpected answer from an LLM, the majority (41 out of 62 participants) said they changed the prompt to try again. A smaller group (12 participants) reported providing feedback to the LLM, while a few (2 participants) said they did nothing (see Figure 6). In the “Other” category, participants described various strategies. Some combine multiple actions, such as changing the prompt, switching to another LLM, or even reverting to traditional search methods like Google or Wikipedia. A few participants mentioned disengaging from the LLM entirely or constructing the solution themselves.
[…]
Participant 39: “Sometimes the LLM hallucinates and puts me in loops. When I realize this, I resolve it myself using my human knowledge and [X] years of experience in the development field”
@@
Figure 6. Actions done after receiving an unexpected answer from LLMs
@@
We asked our participants about the improvements they would like to see that would enhance their experience when using LLMs. They identified several areas that we grouped and sorted in Table 4 based on the number of times they were mentioned. The suggestions mainly align with the engineers’ expectations (See Table 3), though with a variation in the emphasis and distribution.
Regarding the quality of the outcome, the participants emphasised the importance of reducing hallucinations and improving response accuracy. They also commented on the need for enhanced analytical capacity to tackle more complex problems. However, the participants highlighted that while they expect an LLM to be accurate, transparency is more crucial for a less frustrating user experience. The transparency was described by the participants concerned information integrity, particularly specifying the data sources and the confidence level of the LLM’s responses. This also includes stating any assumptions and reasonings by the LLM before providing an answer. This helps practitioners decide whether to rely on the LLM’s suggestions, trust their own judgment, or seek assistance from a colleague instead. Establishing trust in this context is crucial, as it determines the main flow of the interaction, and hence the how the LLM output will be eventually used by the software engineer (Khojah et al., 2024a).
Regarding the answer structure, participants preferred concise and short answers (unless they were prompted otherwise). While in terms of understanding, prompt comprehension was considered essential, with users expecting LLMs to ask clarifying questions when necessary. However, some participants mentioned that this could be a human-related improvement where trainings that are designed specifically to learn to “talk” with the LLM are important to make communication more effective.
Finally, versatility was another improvement that was consistently mentioned among our participants, they indicated the importance of handling more complex tasks and adapting to different user needs. The increased functionality to support a broader range of tasks was also crucial for improving overall user satisfaction. This is especially critical as software engineers researchers and practitioners often work with a variety of tasks that range from learning new concepts and tasks to implementing products and using tools. Furthermore, practitioners particularly stressed the importance LLM performance in terms of increased memory and learning capabilities, allowing LLMs to retain relevant information and build on past interactions with a faster processing time.
Overall, these requirements guide chatbot designers to understand practitioners’ and researchers’ needs and priorities in software engineering. For instance, the next-generation LLMs may need to direct efforts toward transparency rather than performance when aiming for a better user experience.
Table 4. Improvements for Better User Experience
@@
Performance Higher memory utilisation (4)
@@
5. Discussion
In this section, we answer our research questions highlighting the main takeaway per question.
5.1. Frustrations in the Context of LLM Interaction
We identified four main frustration triggers for software engineers using LLMs—accuracy issues, hallucinations, misunderstandings, unmet preferences, and LLMs limitations— pointing to significant challenges with potential long-term repercussions. For example, repeatedly dealing with inaccurate, buggy or non-standard code disrupts workflow, forcing engineers to spend extra time debugging (Fitzgerald et al., 2008) or reworking tasks that an efficient tool should minimise. This is particularly important if the code produced by the LLM is less maintainable (Liu et al., 2024) and can be more frustrating than writing the code with no LLM assistance (Weisz et al., 2022). This extra workload impacts project timelines and creates a deeper frustration that could enforce frustration and emotional strain (Ceaparu et al., 2004; Lazar et al., 2006) as engineers may feel that they have spent longer than planned to reach their goal (e.g., completing a task). Additionally, participants’ frustration resulted from other emotions, including anger and annoyance (see Figure 4).
On Wilcox’s emotion wheel, these frustrations can be linked to various emotions within the frustration spectrum, including anger, annoyance, and confusion. For instance, frustration over inaccurate or faulty code can easily evolve into anger when engineers feel a lack of control over the situation, especially if the tool is supposed to improve efficiency. Similarly, when an LLM produces outputs that deviate from expectations, it can lead to another emotion within the spectrum, for example, annoyance, particularly when the tool fails to meet personal preferences or the engineer’s standards. In addition, misunderstandings or incorrect outputs might lead to confusion as engineers try to reconcile the LLM’s output with their original intent.
Studies on GitHub Copilot (Prather et al., 2023, 2023) showed similar emotions, especially around data privacy concerns, intrusive code suggestions and usability, particularly unnecessary large code suggestions. Eshraghian at al. (Eshraghian et al., 2024) explained that frustration and anger can come from feeling a threat without being able to control it. In the previous example, the threat of leaking confidential data with very little control over it (i.e., to use the LLM, they need to accept the policy) was the trigger for frustration.
Unlike other domains where frustration often stems from performance issues (e.g., system crashes) (Lazar et al., 2006), our participants did not report such frustrations, likely due to recent LLMs being stable and fast (Table 3). Other frustrations in the medical domain arise from the emotionally exhausting work environment along with their dependence on the technology (e.g., to document patient data) (Opoku-Boateng, 2015; Tawfik et al., 2021). In contrast, software engineers can still rely on their expertise or alternative tools, as LLMs are not essential for task completion (Section 4.6).